OpenStudy (anonymous):

The product page numbers on two facing pages of a book is 272. Find the page numbers.

OpenStudy (anonymous):

272 = n(n+1) n^2 + n - 272 = 0 Can you find n?

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

I got -17,16

OpenStudy (anonymous):

Let's call the page number on the left n, and the one on the right n + 1 Then we can express them as the above equation.

OpenStudy (anonymous):

-17 is negative, so reject it 16 is positive, so your two page numbers are 16 and 17

OpenStudy (anonymous):

oh ok...

OpenStudy (anonymous):

1. xy=272 2. x=y-1 , where y is the next page # substitute the value of x on the 1st equation and solve for y (y-1)y= 272 y^2 -y -272 =0 (y-17)(y+16)=0 y=17 or y= -16 (disregard negative value) x = 17-1 = 16